375cmcube of gas has a pressure of 770mmhg. Find it's volume if the pressure is reduced to 750mmhg

1
Expert's answer
2022-05-12T03:55:02-0400

V1=375 cm3=0.000375 m3

P1=770 mmhg=770x133=102658 Pa

P2=750 mmhg=750x133=99991.8 Pa

T,n- constant

V2-?

PV=nRT=C

C=0.000375x102658=38.5

V2=C/P2=38.5/99991.8=0.000385 m3